# Département de mathématiques Université du Québec à Montréal 201, Avenue du Président-Kennedy

### v2

@@ 6 Real-data application
The dataset’s near-symmetry, evident in the histogram, makes the normal distribution a natural first candidate. However, it is rejected at the 5% significance level, with $p$-values of $0.027$ ($T_{n}$) and $0.031$ (LK). The result for the $T_{n}$ test can be interpreted through Figure 4 (left panel), where the point $\smash{[\sqrt{n}C_{n}(\hat{\bm{\theta}}_{n}),\sqrt{n}S_{n}(\hat{\bm{\theta}}_{n})]^{\top}}$ is plotted along with the $95\%$-confidence ellipse for the normal model. The point lies just outside the $95\%$-confidence ellipse, reflecting the fact that the $p$-value is smaller than $0.05$. A univariate analysis of each component shows that the $Z$-scores are $Z(C_{n})=-2.19$ and $Z(S_{n})=1.56$, indicating that the dataset has significantly heavier tails than the normal distribution, since $Z(C_{n})=-2.19$ is negative and more extreme than the left-tail critical value of $Z_{0.025}=-1.96$ under the standard normal distribution. Specifically, the number of observations in the central 50% (the light gray bars in the histogram of Figure 3) is greater than expected under the normal model, while the number of observations in the outer quartiles (the bars with the two darkest shades of gray) is smaller. The value $Z(S_{n})=1.56$, on the other hand, is positive but smaller than $Z_{0.975}=1.96$, suggesting that the distribution is slightly more right-skewed than the normal model, but not significantly so. In summary, the rejection of the normal model is driven by a combination of heavier tails and mild right skewness in the dataset, with the dominant factor being the heavier tails. We observe that the SN model slightly improves the fit relative to the normal distribution by capturing the mild right skewness, with $\hat{\lambda}_{n}=1.539>0$, but it does not improve the fit of the tail thickness. As a result, the corresponding $p$-values, $0.089$ ($T_{n}$) and $0.045$ (LK), lie near conventional significance thresholds, leading to an ambiguous conclusion regarding the null hypothesis.
@@ 7 Summary and outlook
To facilitate widespread adoption, we substantially expanded the scope of both the $T_{n}$ and LK tests by providing the necessary implementation details for $11$ distribution families, yielding $53$ distinct tests that covered most commonly used parametric distributions. We showed that the chi-square $\chi_{2}^{2}$ approximation employed to compute critical values and $p$-values remained very accurate in practice, even for relatively small sample sizes such as $n=30$, for both the $T_{n}$ and LK tests. As a result, these procedures offer broad applicability, comparable to that of classical EDF-based tests, combined with fully plug-and-play implementation, a unique combination to the best of our knowledge. The strong power properties and practical utility of the proposed procedures were illustrated through two simulation studies, a theoretical analysis under local alternatives, and a real-data application involving meteorological forecast errors.
@@
The GitHub repository [7] provides the complete R implementation of the $T_{n}$ and LK goodness-of-fit tests for the 11 families of distributions listed in Tables 1 to 3. It also includes the scripts used to generate all figures, perform the real-data analysis, and produce a numerical validation of all results reported in the paper.
Appendix Appendix B Constants required for the $T_{n}$ and LK test statistics
Table 13 provides a list of constants that must be computed numerically for the evaluation of the $T_{n}$ and LK test statistics. Here, $f_{\mathrm{ga}}(\cdot\,|\,a,b)$ and $\Gamma_{a,b}(\cdot)$ denote the PDF and CDF of a gamma distribution with shape parameter $a$ and scale parameter $b$; $f_{\mathrm{be}}$ and $F_{\mathrm{be}}$ denote the PDF and CDF of a beta distribution, as defined in Table 1; $f_{\mathrm{IG}}$ and $F_{\mathrm{IG}}$ denote the PDF and CDF of an inverse-Gaussian distribution, as defined in Table 1; $\phi$ and $\Phi$ denote the PDF and CDF of the standard normal distribution; and $F_{\mathrm{SN}}$ denotes the CDF of the skew normal distribution, as defined in Table 1.
@@
Appendix Appendix C Supplement for “Omnibus goodness-of-fit tests based on trigonometric moments”
Abstract. This supplement collects the proofs of the results stated in the main document, namely: (i) those supporting the implementation details reported in Tables 2 and 3; and (ii) the proofs of Propositions 4 and 5, which establish the asymptotic behavior of the gamma and EPD tests under local alternatives in Section 5, as well as the corresponding power curves.
C.1 Proof for the $\mathrm{EPD}(\lambda,\mu,\sigma)$ tests in the ML- and MM-based cases
@@
Using the symmetry of $f_{\lambda}(x\,|\,\mu,\sigma)$ with respect to $\mu$ and (C.1), one has
@@
C.2 Proof for the $\mathrm{half\mhyphen EPD}(\lambda,\sigma)$ test
@@
From (C.1), it follows that, for $k\in(-1,\infty)$,
@@
C.3 Proof for the $\mathrm{SN}(\lambda,\mu,\sigma)$ test
@@
C.4 Proof for the $\mathrm{GG}(\lambda,\beta,\rho)$ test
@@
C.5 Proof for the $\mathrm{logistic}(\mu,\sigma)$ test
@@
C.6 Proof for the $\mathrm{Student^{\prime}s}~t(\lambda,\mu,\sigma)$ test
@@
C.7 Proof for the $\mathrm{Gompertz}(\beta,\rho)$ test
@@
C.8 Proof for the $\mathrm{Lomax}(\alpha,\sigma)$ test
@@
C.9 Proof for the $\mathrm{inverse\mhyphen Gaussian}(\mu,\lambda)$ test
@@
C.10 Proof for the $\mathrm{beta}(\alpha,\beta)$ test
@@
C.11 Proof for the $\mathrm{Kumaraswamy}(\alpha,\beta)$ test
@@
C.12 Proof for the $\mathrm{uniform}(a,b)$ test as a limiting case
@@
C.13 Proof of Proposition 4 for the asymptotics of the $\mathrm{gamma}(\lambda,\beta)$ test under local alternatives
@@
C.14 Proof of Proposition 5 for the asymptotics of the $\mathrm{EPD}_{\lambda}(\mu,\sigma)$ test under local alternatives
@@
is given in Section C.1, with $y=(x-\mu_{0})/\sigma_{0}$.
@@
Next, one determines the matrix $S_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})$ for MM estimator. Recall from (14) that
@@
C.15 Power curves for the empirical power study
@@
Figure 6: Power curves of the $T_{n}$, LK, AD, CvM, Ku, and Wa tests for assessing goodness-of-fit to the Student’s $t_{4}$ distribution (with $\lambda_{0}=4$ known and $\mu_{0}$, $\sigma_{0}$ unknown) under $\mathrm{EPD}(0.3\leq\lambda\leq 1.2,\mu,\sigma)$, $\mathrm{EPD}(1.2\leq\lambda\leq 15,\mu,\sigma)$, and $\mathrm{SN}(0\leq\lambda\leq 16,\mu,\sigma)$ alternatives. The nominal significance level is $0.05$, and the sample size is $n=50$.

### v3

@@ 6 Illustrative application to temperature forecast errors
The dataset’s near-symmetry, evident in the histogram, makes the normal distribution a natural first candidate. However, it is rejected at the 5% significance level, with $p$-values of $0.027$ ($T_{n}$) and $0.031$ (LK). The result for the $T_{n}$ test can be interpreted through Figure 4 (left panel), where the point $\smash{[\sqrt{n}C_{n}(\hat{\bm{\theta}}_{n}),\sqrt{n}S_{n}(\hat{\bm{\theta}}_{n})]^{\top}}$ is plotted along with the $95\%$-confidence ellipse for the normal model. The point lies just outside the $95\%$-confidence ellipse, reflecting the fact that the $p$-value is smaller than $0.05$.
@@
A univariate analysis of each component shows that the $Z$-scores are $Z(C_{n})=-2.19$ and $Z(S_{n})=1.56$, indicating that the dataset has significantly heavier tails than the normal distribution, since $Z(C_{n})=-2.19$ is negative and more extreme than the left-tail critical value of $Z_{0.025}=-1.96$ under the standard normal distribution. Specifically, the number of observations in the central 50% (the light gray bars in the histogram of Figure 3) is greater than expected under the normal model, while the number of observations in the outer quartiles (the bars with the two darkest shades of gray) is smaller. The value $Z(S_{n})=1.56$, on the other hand, is positive but smaller than $Z_{0.975}=1.96$, suggesting that the distribution is slightly more right-skewed than the normal model, but not significantly so. In summary, the rejection of the normal model is driven by a combination of heavier tails and mild right skewness in the dataset, with the dominant factor being the heavier tails. We observe that the SN model slightly improves the fit relative to the normal distribution by capturing the mild right skewness, with $\hat{\lambda}_{n}=1.539>0$, but it does not improve the fit of the tail thickness. As a result, the corresponding $p$-values, $0.089$ ($T_{n}$) and $0.045$ (LK), lie near conventional significance thresholds, leading to an ambiguous conclusion regarding the null hypothesis.
@@ 7 Summary and outlook
To facilitate widespread adoption, we substantially expanded the scope of both the $T_{n}$ and LK tests by providing the necessary implementation details for $11$ distribution families, yielding $53$ distinct tests that covered most commonly used parametric distributions. We showed that the chi-square $\chi_{2}^{2}$ approximation employed to compute critical values and $p$-values remained very accurate in practice, even for relatively small sample sizes such as $n=30$, for both the $T_{n}$ and LK tests. As a result, these procedures offer broad applicability, comparable to that of classical EDF-based tests, combined with fully plug-and-play implementation that does not rely on Monte Carlo simulations or pre-tabulated values. To the best of our knowledge, this combination of features is unique. The strong power properties and practical utility of the proposed procedures were illustrated through two simulation studies, a theoretical analysis under local alternatives, and a real-data application involving meteorological forecast errors.
@@
Conflict of interest
The authors declare that they have no conflicts of interest.
@@
The GitHub repository [7] provides the complete R implementation of the $T_{n}$ and LK goodness-of-fit tests for the 11 families of distributions listed in Tables 1 and 2. It also includes the scripts used to generate all figures, perform the real-data analysis, and produce a numerical validation of all results reported in the paper.
Appendix Appendix B Estimators and constants required for the $T_{n}$ and LK test statistics
Table 12 provides the score equations or, where available, explicit estimators for the maximum likelihood estimation of each parameter in the 11 distribution families, together with method-of-moments estimators for the EPD family. Table 13 provides the list of constants $h_{i}(\cdot)$ that must be computed numerically using the R software (for example, using the function cubature::adaptIntegrate) in order to evaluate the $T_{n}$ and LK test statistics. Here, $f_{\mathrm{ga}}(\cdot\,|\,a,b)$ and $\Gamma_{a,b}(\cdot)$ denote the PDF and CDF of a gamma distribution with shape parameter $a$ and scale parameter $b$; $f_{\mathrm{be}}$ and $F_{\mathrm{be}}$ denote the PDF and CDF of a beta distribution, as defined in Table 1; $f_{\mathrm{IG}}$ and $F_{\mathrm{IG}}$ denote the PDF and CDF of an inverse-Gaussian distribution, as defined in Table 1; $\phi$ and $\Phi$ denote the PDF and CDF of the standard normal distribution; and $F_{\mathrm{SN}}$ denotes the CDF of the skew normal distribution, as defined in Table 1.
Table 12: Score equations, or (where available) explicit parameter estimators, for maximum likelihood estimation of each parameter in the 11 distribution families, together with method-of-moments estimators for the EPD family.
@@
Supplementary material
The supplement [8] collects the proofs of the results stated herein, namely: (i) those supporting the implementation details reported in Tables 2 and 12; and (ii) the proofs of Propositions 5.1 and 5.4, which establish the asymptotic behavior of the gamma and EPD tests under local alternatives in Section 5, as well as the corresponding power curves.
Appendix C Proof for the $\mathrm{EPD}(\lambda,\mu,\sigma)$ tests in the ML- and MM-based cases
@@
Using the symmetry of $f_{\lambda}(x\,|\,\mu,\sigma)$ with respect to $\mu$ and (C), one has
@@
Appendix D Proof for the $\mathrm{half\mhyphen EPD}(\lambda,\sigma)$ test
@@
From (C), it follows that, for $k\in(-1,\infty)$,
@@
Appendix E Proof for the $\mathrm{SN}(\lambda,\mu,\sigma)$ test
@@
Appendix F Proof for the $\mathrm{GG}(\lambda,\beta,\rho)$ test
@@
Appendix G Proof for the $\mathrm{logistic}(\mu,\sigma)$ test
@@
Appendix H Proof for the $\mathrm{Student^{\prime}s}~t(\lambda,\mu,\sigma)$ test
@@
Appendix I Proof for the $\mathrm{Gompertz}(\beta,\rho)$ test
@@
Appendix J Proof for the $\mathrm{Lomax}(\alpha,\sigma)$ test
@@
Appendix K Proof for the $\mathrm{inverse\mhyphen Gaussian}(\mu,\lambda)$ test
@@
Appendix L Proof for the $\mathrm{Beta}(\alpha,\beta)$ test
@@
Appendix M Proof for the $\mathrm{Kumaraswamy}(\alpha,\beta)$ test
@@
Appendix N Proof for the $\mathrm{uniform}(a,b)$ test as a limiting case
@@
Appendix O Proof of Proposition 4 for the asymptotics of the $\mathrm{gamma}(\lambda,\beta)$ test under local alternatives
@@
Appendix P Proof of Proposition 5 for the asymptotics of the $\mathrm{EPD}_{\lambda}(\mu,\sigma)$ test under local alternatives
@@
is given in Section C, with $y=(x-\mu_{0})/\sigma_{0}$.
@@
Next, one determines the matrix $S_{\phi,\theta}(\bm{\phi}_{0},\bm{\theta}_{0})$ for $M$ estimator. Recall from (14) that
@@
Appendix Q Power curves for the empirical power study
@@
Figure 6: Power curves of the $T_{n}$, LK, AD, CvM, Ku, and Wa tests for assessing goodness-of-fit to the Student’s $t_{2}$ distribution (with $\lambda_{0}=2$ known and $\mu_{0}$, $\sigma_{0}$ unknown) under $\mathrm{EPD}(0.25\leq\lambda\leq 0.8,\mu,\sigma)$, $\mathrm{EPD}(1.2\leq\lambda\leq 10,\mu,\sigma)$, and $\mathrm{SN}(0\leq\lambda\leq 16,\mu,\sigma)$ alternatives. The nominal significance level is $0.05$, and the sample size is $n=50$.